### v1

# A citation index bridging Hirsch’s $h$ and Egghe’s $g$

Abstract
We propose a citation index $\nu$ (“nu”) and show that it lies between the classical $h$-index and $g$-index. This idea is then generalized to a monotone parametric family $(\nu_{\alpha})$ ($\alpha\geq 0$), whereby $h=\nu_{0}$ and $\nu=\nu_{1}$, while the limiting value $\nu_{\infty}$ is expressed in terms of the maximum citation.
1. Introduction
Hirsch (2005) made a breakthrough in scientometrics by proposing for the first time a simple citation index (commonly referred to as $h$-index), which had an advantage of aggregating the author’s productivity on the basis of both the number of published papers and their quality. Before that, only some extensive summary statistics were used, such as the mean number of citations per paper. Since then, the $h$-index has become a standard metric of authors’ reputation and productivity, for instance routinely taken into consideration in academic appointments and promotions.
Specifically, the $h$-index is defined as the maximum number $h$ of an author’s papers, each cited at least $h$ times (Hirsch, 2005). Therefore, this index only takes into account the fact of a relatively “high” citation of a paper, but the actual number of citations of such a paper is effectively ignored.
To remedy such censoring of larger citations, the alternative citation index (referred to as $g$-index) was proposed by Egghe (2006a), defined as the maximum number $g$ of an author’s most cited papers, such that their total number of citations is at least $g^{2}$. From this definition, it is easy to see that $h\leq g$ (Egghe, 2006b).
These two (by now classical) indexes have attracted a lot of interest and generated ample research into their analytic properties and performance on real datasets, including their estimation in a variety of statistical models of count data (see, e.g., Egghe & Rousseau (2006); Egghe (2006b); Egghe & Rousseau (2008); Yong (2014)). Furthermore, many modifications and alternative variants of the $h$ and $g$ indexes have been proposed, focusing on certain features of the citation profile (see, e.g., Alonso et al. (2009); Guns & Rousseau (2009); Hirsch (2019) and further references therein). An up-to-date online resource SCI2S (2025) is especially useful by keeping track of the state-of-the-art in this domain.
In the present work, we introduce a new citation index $\nu$ (“nu”) aiming to bridge the mathematical definitions of Hirsh’s $h$ and Egghi’s $g$. This idea was first coined in Nuermaimaiti (2023). Namely, we start by observing that the $h$-index can be represented as a sum of certain indicator functions that censor papers to ensure a required minimum of citations. Building on this observation, our $\nu$-index essentially mimics the summative nature of the $h$-index but the new summation explicitly involves the citation numbers of the top papers.
We are then able to show that our $\nu$ is “sandwiched” between $h$ and $g$; more precisely, we prove the two-sided inequalities
$h\leq\nu\leq g^{*},$ (1)
where $g^{*}$ denotes a modified (unconstrained) $g$-index, obtained if we are allowed to add fictitious zeros to the citation vector (Woeginger, 2008b).
On the other hand, a “tempered” version $\bar{\nu}$ of the $\nu$-index, modified so as to be not larger than the number of published papers, satisfies the inequalities
$h\leq\bar{\nu}\leq g.$ (2)
We will finish off by introducing a more general family of citation indexes $(\nu_{\alpha})$, where $\nu_{\alpha}$ is a non-decreasing (integer-valued) function of a real parameter $\alpha\geq 0$. Here, $h=\nu_{0}$ and $\nu=\nu_{1}$, while the limiting value $\nu_{\infty}$ can be expressed in terms of the maximum citation. Concluding remarks will touch upon evaluation of citation indexes in specific frequency models using the concept of limit shape (Nuermaimaiti et al., 2021; Bogachev et al., 2023; Nuermaimaiti, 2023; Yong, 2014).
2. The $h$ and $g$ indexes
[…]
We say that a vector $\bm{x}=(x_{1},\dots,x_{m})$ is dominated by a vector $\bm{y}=(y_{1},\dots,y_{\ell})$ (written as $\bm{x}\preccurlyeq\bm{y}$) if $x_{i}\leq y_{i}$ for all $i\geq 1$; more precisely, if $m\leq\ell$ then $x_{i}\leq y_{i}$ for $i=1,\dots,m$, but if $m>\ell$ then $x_{i}\leq y_{i}$ for $i=1,\dots,\ell$ and $x_{i}=0$ for $i=\ell+1,\dots,m$. Effectively, these two cases imply that we complement the absent components of either $\bm{x}$ or $\bm{y}$ with fictitious zeros to equalize their dimensions, and then the dominance holds component-wise.
The following natural conditions are commonly assumed to be satisfied by any reasonable citation index $c(\bm{x})$ (see Woeginger (2008b); cf. Woeginger (2008a)):
  • (C1)
    If $\bm{x}=\bm{0}$ then $c(\bm{x})=0$.
  • (C2)
    If $\bm{x}=(x_{1},x_{2},\dots,x_{m})$ and $\bm{y}=(x_{1},x_{2},\dots,x_{m},0)$ then $c(\bm{x})=c(\bm{y})$.
  • (C3)
    If $\bm{x}\preccurlyeq\bm{y}$ then $c(\bm{x})\leq c(\bm{y})$.
Let us now recall the above verbal definitions of the $h$ and $g$ indexes and put them into an explicit mathematical formulation.
Starting with the $h$-index, its definition can be expressed as follows,
$h\equiv h(\bm{x})=\max\left\{j\geq 1\colon\!\sum_{i=1}^{m}{\bf 1}_{\{x_{i}\geq j\}}\geq j\right\}\!,$ (3)
where the indicator notation ${\bf 1}_{A_{j}}$ is defined as ${\bf 1}_{A_{j}}=1$ if the condition $A_{j}$ is satisfied, and ${\bf 1}_{A_{j}}=0$ otherwise.
Note that the maximum in (3) is uniquely defined, since the sum on the left-hand side is a decreasing function of $j$, while the right-hand is strictly increasing.
Thus, the sum in (3), for a given $j$, represents the number of papers with at least $j$ citations, while the inequality requires that this number is not smaller than $j$. In particular, noting that the sum of indicators in (3) is not larger than $m$, it follows that the $h$-index is bounded by the number of papers:
$h\leq m.$ (4)
In the degenerate case with $x_{i}\equiv 0$ (i.e., $\bm{x}=\bm{0}$), the inequality in (3) is only satisfied for the value $j=0$, which is excluded from the testing range; thus, the resulting set of suitable $j$’s is empty and, according to the common convention, its maximum is set to be zero: $\max\varnothing=0$; hence
$h(\bm{0})=0,$ (5)
so that property (C1) is automatically satisfied. It is also easy to see that (C2) holds (because $h$ is insensitive to zero citations) and that (C3) is also true.
Next, the definition of the $g$-index can be written mathematically as follows,
$g\equiv g(\bm{x})=\max\left\{1\leq k\leq m\colon\sum_{i=1}^{k}x_{i}\geq k^{2}\right\}\!.$ (6)
Note that if $\bm{x}=\bm{0}$ (i.e., all $x_{i}=0$) then the set under the max-symbol is empty, in which case, by the same convention, we define the maximum as zero. That is to say, the $g$-index for the zero citation vector equals zero:
$g(\bm{0})=0.$ (7)
Also note that, because the testing range of $k$’s in (6) is bounded by $m$, we must have
$g\leq m.$ (8)
It can be shown that the $g$-index is not smaller than the $h$-index of the same author (Egghe, 2006b, Proposition I.2. p. 133),
$h\leq g.$ (9)
Indeed, if the $h$-index has value $h$ then there are $h$ papers with at least $h$ citations each, and therefore with at least $h\times h=h^{2}$ citations in total. Hence, the trial value $k=h$ satisfies the inequality condition in (6), which implies that $g\geq k=h$, as claimed.
Denote $s_{k}=x_{1}+\dots+x_{k}$. According to definition (6), the index $g$ is the largest value of $k$ for which $s_{k}\geq k^{2}$. But it may be unclear whether the inequality $s_{k}\geq k^{2}$ can fail for some $k<g$. Let us show that $s_{k}\geq k^{2}$ for all $k\leq g$.
Lemma 1.
If $s_{k}<k^{2}$ for some $k\geq 1$, then $s_{\ell}<\ell^{2}$ for all $\ell\geq k$.
Proof.
Using that $x_{k}=\min\{x_{1},\dots,x_{k}\}$, we have
$k^{2}>s_{k}=x_{1}+\dots+x_{k}\geq k\mbox{$\>\!$}x_{k},$
which implies that $k>x_{k}\geq x_{k+1}$. Hence,
$s_{k+1}=s_{k}+x_{k+1}<k^{2}+k<(k+1)^{2},$
that is, $s_{k+1}<(k+1)^{2}$. The general claim then follows by induction. ∎
Lemma 2.
If $s_{k}<k^{2}$ then $g<k$. In particular, if $s_{m}<m^{2}$ then $g<m$; otherwise, $g=m$.
Proof.
Readily follows by Lemma 1 and definition (6). ∎
Turning to the verification of the required properties (C1)–(C3) for the $g$-index, we see that (C1) automatically holds due to (7). It is also easy to see that (C3) holds as well. However, the result of Lemma 2 suggests that, a bit surprisingly, property (C2) may fail. For instance, for $\bm{x}=(4)$, we have $g(\bm{x})=1$, but for $\bm{y}=(4,0,0,0)$ (i.e., with three fictitious zeros added), definition (6) yields $g(\bm{y})=2$.
To salvage (C2), and also to amplify the role of top-cited papers, it was suggested (Egghe, 2006b; Woeginger, 2008b) to lift the constraint $k\leq m$ in the definition of the $g$-index (see (6)) by complementing the citation vector $\bm{x}=(x_{1},\dots,x_{m})$ with additional zeros, as if such papers have been published. We denote this version of $g$ by $g^{*}$:
$g^{*}\equiv g^{*}(\bm{x})=\max\left\{k\geq 1\colon\sum_{i=1}^{k}x_{i}\geq k^{2}\right\}\!.$ (10)
This makes no difference if $g\leq m$, because then $g^{*}=g$. However, the case where $g=m$ may be drastically different.
Lemma 3.
Suppose that $s_{m}\geq(m+1)^{2}$. Then $g^{*}=\bigl{\lfloor}\sqrt{s_{m}}\bigr{\rfloor}\geq m+1>m=g$.
Proof.
Adding $g^{*}-m\geq 1$ fictitious zeros to the citation vector $\bm{x}$, we get a new vector $\bm{x}^{\prime}=(x_{1},\dots,x_{m},0,\dots,0)$ of dimension $m+(g^{*}-m)=g^{*}$, resulting in
$\sum_{i=1}^{g^{*}}x^{\prime}_{i}=s_{m}\geq(g^{*})^{2},$
and it follows that $g^{*}=\bigl{\lfloor}\sqrt{s_{m}}\bigr{\rfloor}$. ∎
A striking example illustrating this situation is the case of John Nash (Woeginger, 2008b), with the (rounded) citation vector $\bm{x}=(2000,2000,1500,1000,400,250,100,100)$, for which we get $g=8$ but $g^{*}=85$.
3. An alternative citation index $\nu$
Trying to reconcile the definition of the $h$-index given by formula (3), with the definition of the $g$-index in (6) by taking into account the actual citations of the top papers, we propose a new citation index called the $\nu$-index, defined as the maximum integer $\nu$ such that the total sum of citation counts of papers with at least $\nu$ citations each is not less than $\nu^{2}$. Mathematically, this is expressed as (cf. (3))
$\nu\equiv\nu(\bm{x})=\max\left\{j\geq 1\colon\sum_{i=1}^{m}x_{i}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq j\}}\geq j^{2}\right\}\!.$ (11)
Similarly to (3), the maximum is uniquely defined, noting that the sum in (11) is a decreasing function of $j$, while the right-hand side is strictly increasing. It is also worth pointing out that, unlike $g$ vs. $g^{*}$, the $\nu$-index is insensitive to fictitious zeros.
Simple examples show that the value of $\nu$ may be larger than the total number of papers, in contrast with the $h$ and $g$ indexes. For instance, for $\bm{x}=(9,6,1)$ we get $\nu(\bm{x})=4>3$.
Clearly, this occurs because our definition of $\nu$ gives prominence to few highly cited papers. If unwanted, this can be suppressed by modifying the definition via an explicit constraint $\nu\leq m$:
$\bar{\nu}\equiv\bar{\nu}(\bm{x})=\max\left\{1\leq j\leq m\colon\sum_{i=1}^{m}x_{i}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq j\}}\geq j^{2}\right\}\!.$ (12)
We call $\bar{\nu}$ a tempered $\nu$-index.
Lemma 4.
The indexes $\nu$ and $\bar{\nu}$ satisfy the basic properties (C1)–(C3).
Proof.
Properties (C1) and (C2) are straightforward, since $\nu$ and $\bar{\nu}$ are insensitive to zero values $x_{i}=0$. Monotonicity (C3) is also obvious because the sums in (11) and (12) are monotone increasing in each component $x_{i}$. ∎
The $\nu$-index combines the features of both the $h$-index and the $g$-index. It takes into account citations that are equal to or greater than a minimum threshold value of $\nu$ as in the $h$-index, while also including higher citations as in the $g$-index. This ensures that the $\nu$-index captures the impact of highly cited papers and provides a more balanced picture of their overall scholarly impact. In particular, it may be expected that the $\nu$-index interpolates between $h$ and $g$. The next result supports this conjecture.
Theorem 1.
The citation indexes $h$, $\nu$, $\bar{\nu}$, $g$ and $g^{*}$ are in the following relations:
$h$ $\leq\nu\leq g^{*},$ (13)
$h$ $\leq\bar{\nu}\leq g.$ (14)
Proof.
We only prove inequalities (13); the proof of (14) is similar. First, by definition of $h$ in (3) we can write
$h\leq\sum_{i=1}^{m}{\bf 1}_{\{x_{i}\geq h\}}=\frac{1}{h}\sum_{i=1}^{m}x_{i}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq h\}}.$
Hence,
$\sum_{i=1}^{m}x_{i}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq h\}}\geq h^{2},$
which implies, according to (11), that $\nu\geq h$, as claimed.
Next, let the maximizing sum in (11) be expressed as
$\nu^{2}\leq\sum_{i=1}^{m}x_{i}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq\nu\}}=\sum_{i=1}^{k^{*}}x_{i},$ (15)
thus involving $k^{*}$ terms $x_{i}$, each one satisfying the inequality
$x_{i}\geq\nu\qquad(i=1,\dots,k^{*}).$ (16)
If $k^{*}\leq\nu$ then from (15) we obtain
$\nu^{2}\leq\sum_{i=1}^{k^{*}}x_{i}\leq\sum_{i=1}^{\nu}x_{i},$
and it follows from definition (10) that $g^{*}\geq\nu$. Alternatively, if $k^{*}\geq\nu$ then, using (16), we can write
$\sum_{i=1}^{k^{*}}x_{i}\geq\sum_{i=1}^{\nu}x_{i}\geq\sum_{i=1}^{\nu}\nu=\nu^{2},$
and, as before, it follows that $g^{*}\geq\nu$. ∎
The following Table 1 presents the various citation indexes for a few simple examples.
Table 1: Examples
$\bm{x}=(x_{1},\dots,x_{m})$ $h$ $\bar{\nu}$ $\nu$ $g$ $g^{*}$
$(12,3,1)$ 2 3 3 3 4
$(12,3,1,0)$ 2 3 3 4 4
$(6,3,1,0)$ 2 3 3 3 3
$(4,3,2,2,1)$ 2 2 2 3 3
$(8,1,1)$ 1 2 2 3 3
$(18,18,1,1)$ 2 4 6 4 6
$(20,20,18,6,1,0)$ 4 6 7 6 8
4. Parametric family $(\nu_{\alpha})$
It is quite natural to generalize the definition of the index $\nu$ in (11) by considering different powers. Namely, for $\alpha\geq 0$ we define the $\nu_{\alpha}$-index as
$\nu_{\alpha}\equiv\nu_{\alpha}(\bm{x})=\max\left\{j\geq 1\colon\sum_{i=1}^{m}x_{i}^{\alpha}\mbox{$\;\!$}{\bf 1}_{\{x_{i}\geq j\}}\geq j^{\alpha+1}\right\}\!.$ (17)
Clearly, for $\alpha=0$ and $\alpha=1$ this definition is reduced to (3) and (11), respectively:
$h=\nu_{0},\qquad\nu=\nu_{1}.$ (18)
Like in (3) and (11), the existence and uniqueness of the maximum $\nu_{\alpha}$ is self-evident, noting that the sum is a decreasing function of $j$ while the right-hand side is strictly increasing. It is also straightforward to verify that $\nu_{\alpha}$ satisfies (C1)–(C3). We also observe the monotonicity of the family $(\nu_{\alpha})$.
Lemma 5.
The function $\nu_{\alpha}$ is increasing in $\alpha\geq 0$.
Proof.
Rewrite (17) as
$\nu_{\alpha}=\max\left\{j\geq 1\colon\sum_{i=1}^{m}\left(\frac{x_{i}}{j}\right)^{\mbox{$\>\!\!$}\alpha}{\bf 1}_{\{x_{i}\geq j\}}\geq j\right\}\!,$ (19)
and note that the sum in (19) is monotone increasing in $\alpha$, since $x_{i}/j\geq 1$. ∎
It is also interesting to understand the meaning of the limiting value $\nu_{\infty}$.
Lemma 6.
For a citation vector $\bm{x}=(x_{1},\dots,x_{m})$, denote by $\ell_{1}=\sum_{i=1}^{m}{\bf 1}_{\{x_{i}=x_{1}\}}$ the multiplicity of the top citation number $x_{1}$. Then
$\nu_{\infty}(\bm{x})=\begin{cases}x_{1}-1&\text{if }\,\ell_{1}<x_{1},\\ x_{1}&{\text{otherwise}.}\end{cases}$ (20)
Proof.
Easily follows using (19) by noting that $(x_{1}/j)^{\infty}$ equals $\infty$, $1$ or 00 according as $j<x_{1}$, $j=x_{1}$ or $j>x_{1}$, respectively. ∎
5. Discussion and conclusions
We have introduced some new citation indexes starting with $\nu=\nu_{1}$, and investigated their relations with the classical indexes $h$ and $g$.
As already mentioned, the $h$-index is straightforward and informative, but it is limited by only acknowledging the fact of a high citation but not the actual number of citations. In contrast, the $g$-index is based exclusively on the citations of a few top papers, but ignoring the “footing” of lower-cited papers.
Our synthetic proposal of the $\nu$-index is designed so as to take into account both higher and lower cited papers, which may assess the individual’s productivity in a more fair and balanced way. Indeed, we have seen that the $\nu$-index is in a sense bridging Hirsch’s $h$ and Egghe’s $g$.
Furthermore, the spectrum of the indexes $(\nu_{\alpha})$ provides a flexible toolkit that allows one either to enhance or to inhibit the input from top-cited papers, as required.
Of course, it goes without saying that none of these, or any other indexes known in the literature is perfect and should replace the rest. In fact, a reasonable practical recommendation may be to choose a few indexes to judge someone’s academic achievement, depending on the assessment requirements and also on the specific features of the scientific domain.
In this regard, it may be useful to choose the parameter $\alpha$ in the index $\nu_{\alpha}$ according to certain individual features of the citation vector $\bm{x}$, in the spirit of limit theorems for norms of random vectors (Schlather, 2001; Bogachev, 2006). We will address this issue in our future work.
As another direction of research, it is important to develop methods of evaluation of a preferred citation index under a suitable statistical model of citation frequencies. We advocate (Nuermaimaiti et al., 2021; Bogachev et al., 2023; Nuermaimaiti, 2023) the use of the so-called limit shape of the random citation profile, which enables one to use an appropriate rescaling of the “tail” of the frequency distribution to extract visual information about its behavior. The idea of the limit shape stems from theory of random integer partitions (Vershik, 1996) and was first applied to the citation context by Yong (2014).
To give just one example of utility of the limit shape, consider a power law frequency model $f_{j}=j^{-a}\mbox{$\>\!\!$}/\zeta(a)$ ($j\geq 1$), where $\zeta(a)=\sum_{j\geq 1}j^{-a}$ is the Riemann zeta function ($a>1$). Egghe & Rousseau (2006) proposed the following formula for the $h$-index,
$h\approx m^{1/a},$ (21)
where $m$ is the number of published papers. This expression easily follows upon the replacement of sums emerging in the power law model by the corresponding integrals, which can be explicitly computed. As an alternative, we argue that, even though the power law is scale free (in that $A^{a}f_{\mbox{$\;\!\!$}Ax}=f_{x}$, for any scaling factor $A>0$), the limit shape exists and is given by the formula $y=x^{1-a}/c_{a}$, where $c_{a}=(a-1)\mbox{$\;\!$}\zeta(a)>1$. As a result, the $h$-index can be estimated as follows (Nuermaimaiti, 2023),
$h\approx\left(\frac{m}{c_{a}}\right)^{\mbox{$\>\!\!$}1/a}\!.$ (22)
Estimator (22) is more conservative compared to (21), but most importantly, it retains the number-theoretic flavor of the original discrete distribution (through the term $\zeta(a)$).